# Erasure-tolerance scheme for the surface codes on neutral atom quantum computers

### v3

@@ V.1 Error Models and Assumptions in Our Simulation
Erasure errors occur with probability $p_{\rm e}$ during each round of syndrome measurements, and it remains on the array once an erasure occurs. We suppose that erasure errors can be detected simultaneously before extracting error syndromes without fail or errors for simplicity. We further assume that erasure errors are correctable only if the atom array does not store information of any logical qubit encoded by the surface code, i.e., only incoherent recovery of erased qubits is allowed. This erasure error model causes the accumulation of erasure errors. To simulate these accumulated erasure errors in Stim, we generated $10^{4}$ instances of random erasure occurrence over up to $10d$ rounds of syndrome measurements. Each instance represents the timings and locations of random erasure events obeying an erasure probability. If the instance indicates the erasure occurrence, the completely depolarizing channel
@@
is applied to the erased qubits. We model this process in the simulation by applying a uniformly random Pauli error from $\{I,X,Y,Z\}$.
@@ V.2 Composition of Super Stabilizers
The set of stabilizer generators is composed based on the detected erased qubits. In this study, we formed super stabilizers by Stace’s method. The sets of stabilizer generators are formed into super stabilizers by using a Union-Find data structure, as shown in the pseudocode in Algorithm 1. We can execute this algorithm to generate stabilizer generators for a system with erasures in $O(\log N)$ time steps for $N$ qubits.
@@
As shown in the following section and in Fig. 4, performing multiple rounds of syndrome measurements induces a shorter effective code distance because erasures accumulate along with the rounds. Suppose erasure errors occur continuously between the boundaries, as in Fig. 2, a larger stabilizer generator that connects the boundaries will be composed. The effective code distance of this lattice is $2$, and it does not have sufficient code distance to protect a logical qubit against Pauli errors. Therefore, if erasure errors span from a boundary to the opposite boundary, we grant the logical qubit destroyed in this study.
@@ V.4 Numerical Results on a single array
First, we calculated the threshold of the depolarizing error rate against the logical qubit error rate for each erasure error rate.  Fig. 3 shows the results of sweeping $p_{\textrm{dep}}$ with fixed $p_{\rm e}$. Each data point in Fig. 3 is the average of the logical failure rate of the $10^{4}$ random erasure instances. Each instance simulation runs until $10^{4}$ logical failures are detected or until $10^{5}$ attempts are completed. Hence, each data point is generated by $10^{9}$ attempts at maximum. Each shaded area represents the standard deviation of the sampling. In general, logical failures are unintended flips of logical operators. In this study, disruption of the error-correcting code due to erasure errors is also regarded as a failure.
@@
Fig. 3 shows how the erasure error rate affects the logical failure rate of the surface code. Note that the codes with longer code distances require more repetitions of syndrome measurements, resulting in a higher erasure ratio at the end of the $d$ rounds. As a result, codes with longer code distances degrade more significantly. Due to the faster degradation of longer code distances, the plots for the higher erasure rates in  Fig. 3 do not show the explicit threshold of the depolarizing error rate. In each plot of erasure error rate $p_{\textrm{e}}=1\times 10^{-4}$, $1\times 10^{-3}$ and $1\times 10^{-2}$, there is a threshold-like cross point around $p_{\textrm{dep}}=1\times 10^{-2}$, but not all curves pass through this cross point; it is pseudo-threshold. It suggests that surface codes can reduce the logical failure rate by increasing the code distance up to a certain length, but further increasing the code distance beyond that length does not effectively suppress logical failures. For instance, the graph for the erasure rate of $p_{e}=1\times 10^{-3}$ shows that the surface code can suppress the logical failure rate when the depolarizing error rate is approximately between $3\times 10^{-3}\leq p_{\textrm{dep}}\leq 1\times 10^{-2}$ by increasing the code distance up to $11$. However, the logical failure rate for $d=13$ is worse than for the shorter code distances. Other plots also suggest that long code distances increase the logical failure rate. Thus, low erasure error rates are essential to successfully performing error correction with a longer code distance.
@@
An analytical derivation also supports this result. The derivation indicates that the accumulated-erasure error model does not have a threshold, meaning $p_{\textrm{e, th}}=0$; The ratio of qubits erasing after repeating $d$ times of syndrome measurements is
@@
Figure 3: The logical failure rate after $d$-rounds of syndrome measurements for each erasure error rate $p_{\textrm{e}}$. Each colored line represents a logical failure rate (vertical axis) vs. the depolarizing error rate per gate operation $p_{\textrm{dep}}$ (horizontal axis) for code distances $3$, $5$, $7$, $9$, $11$, and $13$. Each graph has a fixed erasure error rate $p_{\textrm{e}}$. The black dashed line is the break-even line, where the error correction suppresses errors below the depolarizing error rate. Each inset is a zoom-in to the cross-point around $p_{\text{dep}}=10^{-2}$.
@@
Figure 4: This plot shows the decrease in effective code distances with increasing rounds of the syndrome measurements. The horizontal axis $r$ is the repetition number of syndrome measurements normalized by the code distance $d$, i.e., performing the syndrome measurement cycles $dr$ times. Each box plot is obtained from $10^{4}$ circuit samples with $p_{\textrm{e}}=1.0\times 10^{-3}$. The detector error model of every circuit provides the effective code distance as the “shortest graph-like error” in Stim [26]. Each thick black line is the median, and each white triangle is the average of the effective code distance. The whiskers of each box represent the maximum and minimum of the samples. Effective code distances of longer initial code distances decrease rapidly because they are more exposed to accumulated erasure errors due to $d$-repetitions.
@@ VI $k$-Shift Erasure Recovery
To address this problem, we propose an approach that moves the logical qubit from the atom array with accumulated erasure errors to a new atom array free from erased qubits. We call this scheme $k$-shift erasure recovery. This approach can be applied to other types of quantum computers with similar problems, such as qubits getting lost or fidelities getting worse as computation processing and having challenges to fix them while computation is ongoing. Our approach is achieved by two steps, combining the expansion and contraction operations of the surface code through code deformation [21] (Fig. 5):
Figure 5: The process of transferring a logical qubit via code deformation.
@@
This series of steps separates the functionality of executing quantum computation—protecting against state errors and enduring erasure errors—from the functionality of correcting erasure errors. By moving the logical qubit away from the imperfect array, our scheme prevents the logical qubit from suffering effects caused by repairing the atom array, such as noise or time-consuming.
@@
Our protocol performs code deformation [21] to move the logical qubit from $\mathcal{A}_{1}$ to the new perfect array $\mathcal{A}_{2}$ before the logical failure rate becomes too high due to the accumulation of erased qubits. First, initialize the qubits in $\mathcal{A}_{2}$ appropriately to |$\ket{0}$ or $\ket{+}$, depending on the direction of the boundary expansion. Next, perform $d$ rounds of syndrome measurements on $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$ to merge them. Then, the logical qubit is expanded from size $d\times d$ on $\mathcal{A}_{1}$ to size $d\times 2d$ spanning both $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$.
@@
In the same way, erased qubits on the new array $\mathcal{A}_{2}$ will accumulate after repeating error correction. Again, we can make the system tolerant to erasure errors by transferring the logical qubit to the repaired original array $\mathcal{A}_{1}$ before the logical failure rate of $\mathcal{A}_{2}$ becomes too high due to accumulated erasure errors.
The explanation above uses two arrays corresponding to a $2$-shift erasure recovery. This procedure can be generalized to a $k$-shift erasure recovery with $k$ arrays, depending on the time needed to fix erasures on the array and the time until the array can no longer tolerate the accumulation of erasure errors and Pauli errors.
@@
We numerically evaluated the performance of the $k$-shift recovery.  Fig. 6 compares the logical failure rate between performing $3d$ rounds of syndrome measurements on a single array and performing the $2$-shift erasure recovery, which is the minimum setting of the $k$-shift erasure recovery. In the $2$-shift erasure recovery, we perform $d$ rounds of syndrome measurements on $\mathcal{A}_{1}$, then $d$ rounds on the expanded surface code spanning $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, followed by $d$ rounds of syndrome measurements on $\mathcal{A}_{2}$ after the contraction operation, measuring data qubits on $\mathcal{A}_{1}$ and apply feedback to $\mathcal{A}_{2}$. Here, the vertical axis of Fig. 6 represents the logical failure rate of the final state after $3d$ rounds. Each data point is obtained from the average of the logical failure rate of the $10^{4}$ random erasure instances for $3d$ round measurements. As the same as Sec. V.4, each instance simulation runs until $10^{4}$ logical failures are detected or until $10^{5}$ attempts are completed. Hence, each data point is generated by $10^{9}$ attempts at maximum. Each shaded area represents the standard deviation of the sampling. These numerical results show that the $2$-shift recovery improves the logical failure rate compared to the simple repetition of syndrome measurements on a single array. For each code distance (except $d=3$), the $2$-shift recovery achieves suppression of the logical failure rate, whereas performing $3d$ rounds of syndrome measurements on a single array without shifting cannot suppress the logical failure rate at lower depolarizing error rates. This result suggests that $k$-shift recovery provides erasure-tolerance even in its smallest configuration.
@@
Figure 6: Comparison of logical failure rates between conventional syndrome measurements and the $2$-shift erasure recovery for $p_{\textrm{e}}=1.0\times 10^{-3}$. The dotted lines represent the logical failure rate after $3d$ rounds of syndrome measurements on a single array, and the solid lines represent the $2$-shift erasure recovery that takes $3d$ rounds of syndrome measurements, where $d$ is the code distance. The black dashed line is the break-even line, where the error correction suppresses errors below the depolarizing error rate. For each code distance, the $2$-shift recovery performs the same or better logical failure rate, as long as the depolarizing error rate is lower than the threshold around $p_{\textrm{phys}}<10^{-2}$. Each data point is obtained from $10^{9}$ attempts at maximum. The shaded areas represent the standard deviation of the sampling.

### v4

@@ V.1 Error Models and Assumptions in Our Simulation
Erasure errors occur with probability $p_{\rm e}$ before each round of syndrome measurements, and they remains on the array once an erasure occurs. We suppose that erasure errors can be detected simultaneously after extracting error syndromes without fail or errors for simplicity, i.e. the erasure errors affect a single round of syndrome extraction, and we can know the erased qubits after the measurement. We further assume that erasure errors are correctable only if the atom array does not store information of any logical qubit encoded by the surface code, i.e., only incoherent recovery of erased qubits is allowed. This erasure error model causes the accumulation of erasure errors. To simulate these accumulated erasure errors in Stim, we generated $10^{4}$ instances of random erasure occurrence over up to $10d$ rounds of syndrome measurements. Each instance represents the timings and locations of random erasure events obeying an erasure probability. If the instance indicates the erasure occurrence, the completely depolarizing channel
@@
is applied to the erased qubits. We model this process in the simulation by applying a uniformly random Pauli error from $\{I,X,Y,Z\}$ [37, 13].
@@ V.2 Composition of Super Stabilizers
The set of stabilizer generators is composed based on detected erased qubits until the previous measurement rounds, before each syndrome measurement round. In this study, we formed super stabilizers by Stace’s method. The sets of stabilizer generators are formed into super stabilizers by using a Union-Find data structure, as shown in the pseudocode in Algorithm 1. We can execute this algorithm to generate stabilizer generators for a system with erasures in $O(\log N)$ time steps for $N$ qubits.
@@
As shown in the following section and in Fig. 4, performing multiple rounds of syndrome measurements induces a shorter effective code distance because erasures accumulate along with the rounds. Suppose erasure errors occur continuously between the boundaries, as in Fig. 2, a larger stabilizer generator that connects the boundaries will be composed. The effective code distance of this lattice is $2$, and it does not have sufficient code distance to protect a logical qubit against Pauli errors. Therefore, if erasure errors span from a boundary to the opposite boundary, we declare the logical qubit destroyed in this study.
@@ V.4 Numerical Results on a single array
Figure 3: The logical failure rate after $d$-rounds of syndrome measurements for each erasure error rate $p_{\textrm{e}}$. Each colored line represents a logical failure rate (vertical axis) vs. the depolarizing error rate per gate operation $p_{\textrm{dep}}$ (horizontal axis) for code distances 33, 55, 77, 99, 1111, and $13$. Each graph has a fixed erasure error rate $p_{\textrm{e}}$. The black dashed line is the break-even line, where the error correction suppresses errors below the depolarizing error rate. Each inset is a zoom-in to the cross-point around $p_{\text{dep}}=10^{-2}$.
@@
First, we calculated the threshold of the depolarizing error rate against the logical qubit error rate for each erasure error rate.  Fig. 3 shows the results of sweeping $p_{\textrm{dep}}$ with fixed $p_{\rm e}$. Each data point in Fig. 3 is the average of the logical failure rate of the $l=10^{4}$ random erasure instances. Each instance simulation runs until $\xi_{i}=10^{4}$ logical failures are detected or until $\Omega_{i}\leq 10^{5}$ attempts are completed, where $\Omega_{i}$ and $\xi_{i}$ are the number of sampling and the number of logical failure events for each erasure instance, respectively. Hence, each data point is generated by $10^{9}$ attempts at maximum. In general, logical failures are unintended flips of logical operators. In this study, disruption of the error-correcting code due to erasure errors is also regarded as a failure.
Each shaded area in Fig. 3 represents a confidence interval of the maximum-likelihood estimation of logical failure rate $p^{*}={\xi}/{\Omega}$ under the assumption that the data follow a binomial distribution
$f_{\Omega,\xi}(p)=\binom{\Omega}{\xi}\,p^{\xi}(1-p)^{\Omega-\xi},$ (8)
where $\Omega$ and $\xi$ are the total number of sampling and the number of logical failure events, respectively. For this case the shaded interval $p_{\text{low}},\,p_{\text{high}}]$ is defined as the region in which the likelihood falls to at most $1/10^{3}$ of its maximum value, i.e.
$f_{\Omega,\xi}(p_{\text{low}})\;=\;f_{\Omega,\xi}(p_{\text{high}})\;=\;\frac{1}{10^{3}}\;f_{\Omega,\xi}(p^{*}).$ (9)
In this simulation, we performed sampling until $\xi_{i}=10^{4}$ logical failures are detected or until $\Omega_{i}=10^{5}$ attempts for each of $l=10^{4}$ erasure instances. Hence the total number of sampling and the total number of logical failure events are
$\qquad\xi\;=\;\sum_{i=1}^{l}\xi_{i}.$ (10)
Figure 4: This plot shows the decrease in effective code distances with increasing rounds of the syndrome measurements. The horizontal axis $r$ is the repetition number of syndrome measurements normalized by the code distance $d$, i.e., performing the syndrome measurement cycles drdr times. Each box plot is obtained from $10^{4}$ circuit samples with $p_{\textrm{e}}=1.0\times 10^{-3}$. The detector error model of every circuit provides the effective code distance as the “shortest graph-like error” in Stim [30]. Each thick black line is the median, and each white triangle is the average of the effective code distance. The whiskers of each box represent the maximum and minimum of the samples. Effective code distances of longer initial code distances decrease rapidly because they are more exposed to accumulated erasure errors due to $d$-repetitions.
@@
Fig. 3 shows how the erasure error rate affects the logical failure rate of the surface code. Note that the codes with longer code distances require more repetitions of syndrome measurements, resulting in a higher erasure ratio at the end of the $d$ rounds. As a result, codes with longer code distances degrade more significantly. Due to the faster degradation of longer code distances, the plots for the higher erasure rates in  Fig. 3 do not show the explicit threshold of the depolarizing error rate. In each plot of erasure error rate $p_{\textrm{e}}=1\times 10^{-4}$, $1\times 10^{-3}$ and $1\times 10^{-2}$, there is a threshold-like cross point around $p_{\textrm{dep}}=1\times 10^{-2}$, but not all curves pass through this cross point; it is a pseudo-threshold. It suggests that surface codes can reduce the logical failure rate by increasing the code distance up to a certain length, but further increasing the code distance beyond that length does not effectively suppress logical failures. For instance, the graph for the erasure rate of $p_{e}=1\times 10^{-3}$ shows that the surface code can suppress the logical failure rate when the depolarizing error rate is approximately between $3\times 10^{-3}\leq p_{\textrm{dep}}\leq 1\times 10^{-2}$ by increasing the code distance up to $11$. However, the logical failure rate for $d=13$ is worse than for the shorter code distances. Other plots also suggest that long code distances increase the logical failure rate. Thus, low erasure error rates are essential to successfully performing error correction with a longer code distance.
@@
An analytical derivation also supports this result. The derivation indicates that the accumulated-erasure error model does not have a threshold, meaning $p_{\textrm{e, th}}=0$; the ratio of qubits erased after repeating $d$ times of syndrome measurements is
@@
(a) Transferring a logical qubit
(b) Timeline of the 3-shift erasure recovery
Figure 5: (a) The process of transferring a logical qubit via code deformation. (b) The timeline of the 3-shift erasure recovery process. The horizontal axis represents the indexing of three arrays used to store the information of the logical qubit. The orange blocks indicate the location of the logical qubit, while the shaded blue blocks denote the measurement of data qubits performed to contract the extended logical qubit. The red blocks indicate the relocation of an array to the position where it will be utilized in the next step, depending on the geometric configuration of the tweezer arrays. Initially, the logical qubit is allocated to $\mathcal{A}_{1}$. Subsequently, the logical qubit is expanded to span $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$. Following this, the logical qubit is contracted to $\mathcal{A}_{2}$ by measuring the data qubits on $\mathcal{A}_{1}$. Finally, $\mathcal{A}_{1}$ is shifted to the adjacent to $\mathcal{A}_{3}$, as $\mathcal{A}_{1}$ will receive the logical qubit from $\mathcal{A}_{3}$ in the subsequent step.
@@ VI $k$-Shift Erasure Recovery
To address this problem, we propose an approach that moves the logical qubit from the atom array with accumulated erasure errors to a new atom array free from erased qubits. We call this scheme $k$-shift erasure recovery. This approach can be applied to other types of quantum computers with similar problems, such as qubits getting lost or fidelities decreasing as computation processes in a manner difficult to fix while computation is ongoing. Our approach is achieved by two steps, combining the expansion and contraction operations of the surface code through code deformation [25] (Fig. 5(a)):
@@
This series of steps separates the functionality of executing quantum computation—protecting against state errors and enduring erasure errors—from the functionality of correcting erasure errors. By moving the logical qubit away from the imperfect array, our scheme prevents the logical qubit from suffering effects caused by repairing the atom array, such as noise or increased errors during the induced idling time.
@@
[…]
@@
In the same way, erased qubits on the new array $\mathcal{A}_{2}$ will accumulate after repeating error correction. Again, we can make the system tolerant to erasure errors by transferring the logical qubit to another array $\mathcal{A}_{3}$ before the logical failure rate of $\mathcal{A}_{2}$ becomes too high due to accumulated erasure errors. Repeating transferring the logical qubit (k1)(k-1) times secures a sufficient time of the restoration of the array $\mathcal{A}_{1}$. Thus, the logical qubit on $\mathcal{A}_{k}$ can be transferred onto $\mathcal{A}_{1}$. Repetition of this cycle prevents logical qubit from a logical Pauli error and erasure errors.
The number of the redundant arrays $k$ is decided depending on the time needed to fix erasures on the array and the time until the array can no longer tolerate the accumulation of erasure errors and Pauli errors. In this protocol, the number of arrays $k$ contributes to redundancy in the loss of information at the code-patch level. Note that this redundancy is principally enough to use $k=2$ arrays if resorting of erased atoms is quickly accomplished. However, if enough $k$ arrays are not prepared, this redundancy cannot be utilized. It is analogous to the bus error in classical computer architecture. This events will cause the information loss of a logical qubit memory. Therefore, in order to secure sufficient time to restore the atom arrays, this scheme introduces redundancy by increasing the number $k$ of arrays.
@@
To evaluate the performance of the $k$-shift recovery, the most important point is that transferring logical qubit with accumulated erasure errors effectively works. We numerically evaluated the effectiveness of the transferring logical qubits with accumulated erasure errors, i.e. $2$-shift erasure recovery. Fig. 6 compares the logical failure rate between performing $3d$ rounds of syndrome measurements on a single array and performing the $2$-shift erasure recovery, which is the minimum setting of the $k$-shift erasure recovery. In the $2$-shift erasure recovery, we perform $d$ rounds of syndrome measurements on $\mathcal{A}_{1}$, then $d$ rounds on the expanded surface code spanning $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, followed by $d$ rounds of syndrome measurements on $\mathcal{A}_{2}$ after the contraction operation and measuring data qubits on $\mathcal{A}_{1}$ with Pauli Z-basis. Therefore, the ”2-shift” data points are generated by simulations of 3 rounds in total. And ”single array” data points are generated with 3 rounds of single arrays for comparison. Pauli errors on the circuit are decoded by Pymatching with the detector error model of the above circuit. Here, the vertical axis of Fig. 6 represents the logical failure rate of measuring the final state on logical $Z$ basis after $3d$ rounds. Each data point is obtained from the average of the logical failure rate of the $10^{4}$ random erasure instances for $3d$ round measurements. As the same as Sec. V.4, each instance simulation runs until $10^{4}$ logical failures are detected or until $10^{5}$ attempts are completed. Hence, each data point is generated by $10^{9}$ attempts at maximum. Each shaded area represents the confidence interval of the maximum-likelihood estimation of logical failure rate in the same way shown in Sec. V.4. These numerical results show that the $2$-shift recovery improves the logical failure rate compared to the simple repetition of syndrome measurements on a single array. For each code distance (except $d=3$), the $2$-shift recovery achieves suppression of the logical failure rate, whereas performing $3d$ rounds of syndrome measurements on a single array without shifting cannot suppress the logical failure rate at lower depolarizing error rates. Shifting the logical qubit to the new array effectively resets the number of erased qubits that a logical qubit holds to zero. Thus, the continuous accumulation of erased qubits is prevented. Therefore, $2$-shift erasure recovery can more effectively suppress the logical failure rate than single array cases. This result suggests that $k$-shift recovery provides erasure-tolerance even in its smallest configuration since the proposed scheme works between patch $N$ and patch $N+1$ as well. Even if we conduct end-to-end simulations of such as $k=10$, failure rate per stabilizer round is considered the same, since erasure errors are refreshed at the beginning of each round.
@@
Figure 6: Comparison of logical failure rates between conventional syndrome measurements and the $2$-shift erasure recovery for $p_{\textrm{e}}=1.0\times 10^{-3}$. The dotted lines represent the logical failure rate after $3d$ rounds of syndrome measurements on a single array, and the solid lines represent the $2$-shift erasure recovery that takes $3d$ rounds of syndrome measurements, where $d$ is the code distance. The black dashed line is the break-even line, where the error correction suppresses errors below the depolarizing error rate. For each code distance, the $2$-shift recovery performs the same or better logical failure rate, as long as the depolarizing error rate is lower than the threshold around $p_{\textrm{phys}}<10^{-2}$. Each data point is obtained from $10^{9}$ attempts at maximum. Each shaded area represents the confidence interval of the sampling defined in (9). Note that the cross point of solid lines is a pseudo-threshold as the same as the single-patch simulation in Fig. 3.
We also calculate $k_{\mathrm{req}}$, the number of arrays required for a logical qubit to pass before returning to its original array so that a defective array has sufficient time to be restored to a perfect one. The time scale to measure a physical qubit by imaging the fluorescence photons from a single atom is milliseconds [39], while the time scale for gate operations is typically microseconds [40]. Thus, the measurement on physical qubits is the most time-consuming operation in a syndrome measurement. We assume that measuring the syndromes of all stabilizer generators in a single round takes $10$ ms, including erasure detection. For the code with distance $d$, repeating $r$ rounds of syndrome measurements will consume $T_{\textrm{meas}}=10rd$ ms. On the other hand, sorting to a two-dimensional defect-free array also takes hundreds of milliseconds. Suppose we have a square lattice array of $N=(2d-1)^{2}+2d$ atoms if the array is perfect. In Ref. [6], the time scale of sorting is estimated to be $0.28N^{1.4}$ ms. Therefore, the time scale to sort to a perfect array is $T_{\textrm{sort}}\propto O(d^{2.8})$.
Figure 7: The dependency of $T_{\textrm{sort}}/T_{\textrm{meas}}$ for code distance on each number of repetition rounds $r$.This ratio suggests the lower bound of $k$. This figure shows that the code with longer code distance requires more atom arrays.
$k_{\mathrm{req}}$ is lower bounded by the ratio of $T_{\textrm{meas}}$ and $T_{\textrm{sort}}$ since larger $k$ in this scheme allow longer time to recover erasures on the array. Fig. 7 shows the ratio of $T_{\textrm{sort}}/T_{\textrm{meas}}$. Here, $r$ is the number of repeated rounds of error correction cycles where each cycle consists of $d$ syndrome measurements performed on the logical qubit in a given array before transferring to the next array. For example of $r=2$, $d$ rounds of syndrome measurements are performed on a single array, then $d$ rounds of syndrome measurements are performed on a $d\times 2d$ array before contracting data qubits on the array containing erased qubits. The figure suggests that a longer code distance requires that a logical qubit is transferred over more arrays. When d=9,r=2d=9,\ r=2 which is the largest setting in the numerical simulation in Fig. 6, $T_{\textrm{sort}}/T_{\textrm{meas}}\simeq 4.72$. Therefore, $k=5$ arrays are needed to protect a logical qubit. Conversely, $k_{\mathrm{req}}$ decreases if the erasure rate allows more error correction cycles at each array. Suppose an experimentalist sets the threshold of erased qubit ratio
$R_{\mathrm{th}}=1-(1-p_{\mathrm{e}})^{\alpha rd}.$ (12)
$k_{\mathrm{req}}$ is lower bounded by
$k_{\mathrm{req}}\geq T_{\textrm{sort}}/T_{\textrm{meas}}\propto\left[(2d-1)^{2}+2d\right]^{1.4}\frac{\log(1-p_{\mathrm{e}})}{\log(1-R_{\mathrm{th}})}.$ (13)
Thus, $k_{\mathrm{req}}$ is minimized with respect to the code distance and the erasure error rate.